OpenStudy (anonymous):

the average of a and b is 5, and the average f c,d, and 10 is 24. what is the average of a, b, c and d?

OpenStudy (cwrw238):

by definition of the average a + b = 10 (since the average = 5) also c + d + 10 = 3 * 24 = 72 c + d = 72-10 = 62 so a + b + c + d = 62 + 10 = 72 average of a,b,c and d = 72/4 = 18